A number between 300 and 400. If it is divided by 2, the remainder is 1. If it is divided by 4,6, or 8, the remainder is 3. If it is devied by 10, the remainder is 5. If t divided by 3,5,7, or 9, the remainder is zero. What is the number?
Answer by ankor@dixie-net.com(22740)

A number between 300 and 400. If it is divided by 2, the remainder is 1. If it is
divided by 4,6, or 8, the remainder is 3. If it is divided by 10, the remainder
is 5. If t divided by 3,5,7, or 9, the remainder is zero. What is the number?
:
" If it is divided by 10, the remainder is 5." that means the number ends in 5
:
Using the statement; "If t divided by 3,5,7, or 9, the remainder is zero."
5 * 7 * 9 = 315
:
What is the number? 315
:
Check solution in the other statements
If it is divided by 2, the remainder is 1: It's an odd number
If it is divided by 4,6, or 8, the remainder is 3.
315/4 = 78 r3
315/6 = 52 r3
315/8 = 39 r3